A lead bullet moves horizontally with velocity of 130 m-s-1 and embedded into a cement wall after collision. If the specific heat capacity of lead =130J. Kg-1.°c-1 and all the heat produced is absorbed by the bullet what is the increase in temperature of the bullet?


1
Expert's answer
2021-03-08T08:26:09-0500

Kinetic energy of the bullet


"E_K=\frac12 mv^2"


fully converts into heat:


"E_K=E_H=cm\Delta t,\\\space\\
\Delta t=\frac{E_K}{cm}=\frac{v^2}{2c}=65º\text{C}."